40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11198     Entry time: Fri Apr 3 10:46:32 2015
Author: Steve 
Type: Update 
Category: SUS 
Subject: OSEM sensor oscillation 

ITMX, ETMY, BS and SRM are oscillating ?

Attachment 1: sensors1sec.png  86 kB  | Show | Show all
ELOG V3.1.3-